# CAD-Unet: A Capsule Network-Enhanced Unet Architecture for Accurate Segmentation of COVID-19 Lung Infections from CT Images

### v1

@@ 4.4 Experimental results
Table 5: Comparative experimental results of multi-label segmentation task on dataset 1 and dataset 2.
@@
Fig. 6: Visual comparison of a segmentation model trained with different segmentation architectures for multi-classes (No-infection, GGO and CON) COVID-19 infection segmentation using dataset 1 and dataset 2. GGO is presented by the green color and CON by the red color.
@@ 4.5 Ablation study
To further validate the efficacy of CAD-Unet and its individual components, ablation experiments are conducted on four datasets for two tasks. Binary segmentation results on dataset 2, dataset 3, and dataset 4 are summarized in Table 6, while multi-class segmentation results on dataset 1 and dataset 2 are presented in Table 7. These experiments evaluate the effects of the capsule layer, dual decoders, and attention gate.
@@
In the binary segmentation task, we initially examine the enhancement of the CAD-Unet module on the baseline model (Unet). As shown in No.1 and No.2 of Table 6, incorporating the capsule module leads to improvements in F1 score by 2.80, 1.40, and 0.65 on the three datasets, with varying degrees of enhancement in other metrics. Subsequently, the impact of the capsule, dual decoders, and attention gate modules on the binary segmentation capability of the proposed CAD-Unet is investigated, from No.3 to No.6 as depicted in Table 6. Disabling the capsule results in a decrease in F1 score by 11.44, 1.51, and 1.01 on the three datasets, which indicates the critical role of the capsule layer in capturing hierarchical spatial relationships and enhancing feature representation, significantly contributing to the accuracy of the model. Following the deactivation of the attention gate, F1 scores decrease by 6.74, 0.39, and 0.24 on the three datasets, while deactivating the dual decoders lead to even larger decreases of 15.51, 2.51, and 0.34. These findings suggest that the attention gate and dual decoders in CAD-Unet also bring significant improvements to segmentation performance in binary segmentation tasks. The introduction of dual decoders helps the network focus on the lung region, while the attention gate design enables the model to concentrate more on crucial spatial information, thereby enhancing the recognition of subtle features.
@@
In the multi-class segmentation task, similar ablation experiments are conducted. As shown in Table 7, akin to the results in the binary segmentation task, the CAD-Unet module demonstrates a noticeable performance improvement over the baseline model (Unet). Incorporating the capsule into the baseline model results in F1 score improvements of 2.09 and 23.93 percentage points for GOO and CON classifications, respectively. Disabling the capsule in CAD-Unet leads to decreases of 3.2 and 3.09 in F1 score for GOO and CON classifications. This indicates that the capsule module plays a crucial role in capturing hierarchical relationships between different categories, enhancing the model’s ability to handle multi-class segmentation tasks. Following this, the effects of the attention gate and dual decoders on multi-class segmentation capability are validated. Disabling the attention gate in CAD-Unet results in F1 score decreases of 2.62 and 3.24 for GOO and CON classifications, while disabling the dual decoders leads to F1 score decreases of 0.87 and 3.72 for GOO and CON classifications. This suggests that the attention gate and dual decoders modules also play positive roles in enhancing model performance in multi-class segmentation tasks.
@@
In summary, through in-depth analysis of ablation experiments, we conclude that the capsule module, dual decoders, and attention gate in CAD-Unet are indispensable components, collectively contributing to the outstanding performance of the model in medical image segmentation tasks. These experiments confirm the effectiveness of our proposed architecture and provide valuable insights for future research in deep learning for medical image segmentation. In future, we will continue to optimize these components, expand their applicability, and explore other potential methods to enhance model performance.
No. Architecture Ablation Dataset 2 Dataset 3 dataset 4
Cap AG DD F1-S IoU Rec F1-S IoU Rec F1-S IoU Rec
1 Unet(baseline) $\times$ $\times$ $\times$ 67.37 50.8 58.09 74.70 59.62 75.81 79.22 66.56 84.22
2 Cap $\checkmark$ $\times$ $\times$ 70.17 54.05 59.46 76.10 61.42 81.60 79.89 66.52 86.52
3 AG&DD $\times$ $\checkmark$ $\checkmark$ 69.97 53.81 57.44 76.62 62.11 81.15 79.81 66.40 85.51
4 Cap&DD $\checkmark$ $\times$ $\checkmark$ 74.87 59.84 65.53 77.74 63.59 80.48 80.66 67.59 86.12
5 Cap&AG $\checkmark$ $\checkmark$ $\times$ 66.10 49.36 52.23 75.62 60.80 80.11 80.56 67.45 85.07
6 Cap&AG&DD 81.61 68.93 82.79 78.13 64.11 86.01 80.90 67.93 87.19
No. Architecture Ablation GGO CON
Cap AG DD F1-S IoU Rec F1-S IoU Rec
1 Unet(baseline) $\times$ $\times$ $\times$ 65.81 50.13 49.06 31.35 15.45 19.26
2 Cap $\checkmark$ $\times$ $\times$ 67.90 51.45 59.40 55.28 38.22 6131
3 AG&DD $\times$ $\checkmark$ $\checkmark$ 67.51 51.09 58.99 56.52 39.41 59.47
4 Cap&DD $\checkmark$ $\times$ $\checkmark$ 68.09 51.67 58.97 56.37 39.25 62.86
5 Cap&AG $\checkmark$ $\checkmark$ $\times$ 69.84 53.67 63.53 55.89 38.80 58.47
6 Cap&AG&DD $\checkmark$ $\checkmark$ $\checkmark$ 70.71 54.69 63.34 59.61 42.47 64.59
Table 7: The results of ablation experiments for multi-segmentation tasks. The influence of three modules, namely Cap, AG, and DD, on the overall architectural performance in multi-valued segmentation is verified on the combined dataset of dataset 1 and dataset 2.
@@ 5 Discussion
To visually demonstrate the efficacy of our proposed model in addressing both binary and multi-class segmentation tasks for COVID-19 lung CT lesion identification, we conducted a qualitative examination of the segmentation outcomes for both scenarios. Figure 5 provides a visual representation of partially predicted masks alongside the ground truth (GT) for the binary segmentation task, derived from three distinct datasets. The proposed model is compared with Unet++, nCoVSeg, SCOAT-Net, and PDAtt-Unet models. The results clearly demonstrate the superior performance of the predicted masks by our model, surpassing the other four models in terms of accurately encompassing the entire lesion region and exhibiting more precise and well-defined segmentation boundaries.
This underscores the capability of the model in effectively addressing binary segmentation challenges associated with COVID-19 lung lesions. The predicted masks produced by our model consistently demonstrate a propensity to encapsulate intricate lesion shapes while displaying sharper and more precise boundaries compared to the alternative models.
Figure 6 illustrates the visualization of partially predicted masks and GT for the multi-class segmentation task from two datasets, using our proposed model along with Unet++, SCOAT-Net, Anamnet, and D-TrAttUnet models. Comparing the predicted masks and GT of other models, the first row of images shows that the proposed model captures more complete CON class masks. The second row demonstrates the superior segmentation performance of our model on small target regions. The third row indicates that our model obtains more accurate target shape information, while the fourth row shows that we achieve more complete target segmentation masks in the GOO class.
@@ 6 Conclusions
The outbreak of COVID-19 has presented significant challenges to global health. Lung CT imaging plays a crucial role in the early diagnosis and treatment of COVID-19 patients. This paper aims to develop an efficient and accurate lung CT lesion segmentation method to assist clinicians in quickly obtaining lesion information for more precise diagnosis and treatment. Our proposed CAD-Unet architecture combines the Unet encoder pathway and capsule encoder pathway in parallel to extract richer local, global, and long-range dependent features. Furthermore, to address the limitations of the BCE loss function in segmenting infection edges, we have incorporated an edge loss component.
To assess the effectiveness of the proposed method, we conducted evaluations on two tasks: binary segmentation and multi-class segmentation. In addition, we compared our method with three baselines and five state-of-the-art architectures for COVID-19 segmentation. The experimental results demonstrate that our method outperforms all the mentioned architectures in common evaluation metrics for both binary and multi-class segmentation tasks. Furthermore, we performed ablation experiments to validate the effectiveness of the components in our proposed model. In future work, we will explore the potential of combining capsule networks with traditional segmentation models and applying them to a wider range of medical image segmentation tasks.

### v2

@@ 4.4.2 Multi-class segmentation task
Table 7: Comparative experimental results of multi-class segmentation task on dataset 1 and dataset 2.
Model Dataset 2 Dataset 3 Dataset 4
Unet 0.009 0.009 0.009
UNet++ 0.009 0.047 0.009
Att-Unet 0.009 0.047 0.009
InfNet 0.009 0.009 0.009
SCOATNet 0.009 0.028 0.009
nCoVSegNet 0.047 0.009 0.009
PDEAtt-Unet 0.009 0.009 0.009
D-TrAttUnet 0.009 0.009 0.009
Unet(Hybrid Loss) 0.009 0.009 0.009
UNet++(Hybrid Loss) 0.009 0.009 0.009
Att-Unet(Hybrid Loss) 0.009 0.028 0.009
SCOATNet(Hybrid Loss) 0.009 0.009 0.009
nCoVSegNet(Hybrid Loss) 0.047 0.009 0.009
Table 8: P-values for binary segmentation across datasets.
Model GGO CON
Unet 0.009 0.009
Att-Unet 0.047 0.009
Unet++ 0.009 0.009
CopleNet 0.009 0.009
AnamNet 0.009 0.009
SCOATNET 0.009 0.009
D-TrAttUnet 0.32 0.045
Table 9: P-values for multi-class segmentation.
Architecture Ablation Dataset 2 Dataset 3
CAP AG DD F1 score IoU Recall F1 score IoU Recall
ResUnet(baseline) $\times$ $\times$ $\times$ 53.14±8.45 36.62±7.51 43.84±11.08 74.99±1.76 75.99±1.76 78.40±3.94
AttUnet(baseline) $\times$ $\checkmark$ $\times$ 63.63±8.74 46.77±6.02 48.96±7.37 76.25±1.05 61.09±0.99 71.86±1.55
Cap $\checkmark$ $\times$ $\times$ 61.74±7.00 45.02±7.27 51.74±10.63 75.39±0.61 60.31±0.79 77.53±3.03
AG&DD $\times$ $\checkmark$ $\checkmark$ 71.06±5.90 55.41±6.69 63.68±5.35 77.34±0.63 63.05±0.85 79.60±1.48
Cap&DD $\checkmark$ $\times$ $\checkmark$ 73.12±3.80 57.77±4.72 64.32±4.74 77.24±1.29 62.94±1.72 81.23±3.32
Cap&AG $\checkmark$ $\checkmark$ $\times$ 64.97±7.13 48.50±7.32 54.45±6.65 74.93±0.67 59.91±0.86 80.80±2.62
Cap&AG&DD $\checkmark$ $\checkmark$ $\checkmark$ 79.82±1.89 64.46±2.64 78.73±3.81 77.97±0.72 63.90±0.98 82.98±1.38
Architecture Ablation Dataset 4
CAP AG DD F1 score IoU Recall
ResUnet(baseline) $\times$ $\times$ $\times$ 76.79±0.41 65.00±0.56 83.23±0.87
AttUnet(baseline) $\times$ $\checkmark$ $\times$ 77.20±0.57 66.95±0.80 77.18±1.01
Cap $\checkmark$ $\times$ $\times$ 80.49±0.32 67.36±0.45 86.17±0.53
AG&DD $\times$ $\checkmark$ $\checkmark$ 80.43±0.35 66.00±0.48 85.62±1.97
Cap&DD $\checkmark$ $\times$ $\checkmark$ 80.51±0.40 65.89±0.55 85.25±1.77
Cap&AG $\checkmark$ $\checkmark$ $\times$ 80.21±0.17 65.58±0.23 84.28±1.08
Cap&AG&DD $\checkmark$ $\checkmark$ $\checkmark$ 80.83±0.20 67.83±0.28 86.17±1.72
Architecture Ablation GGO CON
CAP AG DD F1 score IoU Recall F1 score IoU Recall
ResUnet(baseline) $\times$ $\times$ $\times$ 68.41±1.89 52.02±1.41 61.27±2.42 54.39±2.05 37.89±1.95 59.61±5.51
AttUnet(baseline) $\times$ $\checkmark$ $\times$ 69.66±0.88 53.46±1.04 62.05±1.33 54.61±1.62 37.58±1.53 60.18±2.62
Cap $\checkmark$ $\times$ $\times$ 67.90±2.55 51.45±2.88 59.40±4.73 55.28±1.95 38.22±1.86 61.31±1.90
AG&DD $\times$ $\checkmark$ $\checkmark$ 67.51±3.94 51.09±4.44 58.99±5.87 56.52±1.82 39.41±1.79 59.47±4.59
Cap&DD $\checkmark$ $\times$ $\checkmark$ 68.09±2.53 51.67±2.86 58.97±4.58 56.37±0.58 39.25±0.56 62.86±2.65
Cap&AG $\checkmark$ $\checkmark$ $\times$ 69.84±1.43 53.67±1.66 63.53±1.34 55.89±1.50 38.80±1.44 58.47±0.80
Cap&AG&DD $\checkmark$ $\checkmark$ $\checkmark$ 70.71±0.44 54.69±0.52 63.34±0.98 59.61±0.74 42.47±0.75 64.59±2.14
Table 11: The results of ablation experiments for multi-class segmentation tasks. The influence of three modules, namely Cap, AG, and DD, on the overall architectural performance in multi-class segmentation is verified on the combined datasets of dataset 1 and dataset 2.
@@
Fig. 6: Visual comparison of a segmentation model trained with different segmentation architectures for multi-class (No-infection, GGO and CON) COVID-19 infection segmentation using dataset 1 and dataset 2. GGO is represented in green and CON in red.
4.5 Statistical tests
To test the statistical significance of the proposed model’s performance, we conduct statistical tests on the F1 scores for both tasks. Given the non-normal distribution of performance metrics typical in deep learning, we adopt the Wilcoxon rank-sum test. The F1 scores of the proposed model and compared models are treated as independent samples, and p-values are computed for each pairwise comparison. A p-value $<$ 0.05 indicates that the observed differences are unlikely to be due to chance, providing evidence against the null hypothesis and supporting the superiority of the proposed model. The p-values for both tasks are summarized in Tables 8 and 9.
For the binary segmentation task, across all three datasets, the p-values between the proposed model and all compared models are below 0.05, indicating that the performance improvements of the proposed model are statistically significant.
For the multi-class segmentation task, the proposed model demonstrates significant improvements over most baseline models.Notably, in the GGO classification, the p-value of 0.32 when compared to the D-TrAttUnet model indicates no statistically significant difference between the two models. However, in the CON classification, the p-value of 0.045 suggests a significant performance difference. Additionally, the proposed model exhibits a lower standard deviation in the GGO classification, highlighting its stability. Given the sensitivity of medical image segmentation tasks to stability, these results suggest that the proposed model remains a strong candidate despite the lack of statistical significance in certain cases.
@@ 4.6 Ablation study
To evaluate the effectiveness of CAD-Unet and its individual components, ablation experiments are conducted across four datasets and two tasks. Referring to the viewpoints in [8], and to avoid baseline-related pitfalls, we employ ResUnet, a U-Net with residual connections, and AttUnet as baseline models. The results for binary segmentation on datasets 2, 3, and 4 are summarized in Table 10, while those for multi-class segmentation on datasets 1 and 2 are presented in Table 11. These experiments assess the contributions of the capsule layer, dual decoders, and attention gate.
@@
In the binary segmentation task, we first evaluate the impact of the CAD-Unet module on the baseline ResUnet model. As shown in Table 10, incorporating the attention gate and capsule module individually leads to improvements in F1 scores by 10.49, 1.26, 0.41 and 8.6, 0.4, 3.7 for datasets 2, 3, 4, respectively. The enhancements in other metrics also vary, indicating that both the capsule module and attention gate contribute to performance improvement. Next, we examine the combined effects of the capsule, dual decoders, and attention gate on binary segmentation performance. Disabling the capsule results in F1 score reductions of 8.76, 0.63, and 0.4 on the three datasets, highlighting the importance of the capsule layer in capturing hierarchical spatial relationships and enhancing feature representation. The deactivation of the attention gate causes F1 score reductions of 6.7, 0.73, and 0.32, while disabling the dual decoders leads to even larger reductions of 14.85, 3.04, and 0.62. These results suggest that the attention gate and dual decoders also significantly contribute to performance improvements in binary segmentation tasks. The dual decoders help the network focus on the lung region, while the attention gate design enables the model to concentrate on critical spatial information, enhancing the recognition of subtle features.
@@
In the multi-class segmentation task, similar ablation experiments are conducted. As shown in Table 11, incorporating the attention gate module into the baseline model results in F1 score improvements of 1.25 and 0.22 percentage points for GGO and CON classifications, respectively. Incorporating the capsule module into the ResUnet model results in an F1 score decrease of 0.51 for GGO classification but an improvement of 0.89 for CON classification, which indicates that the standalone capsule module performs poorly for GGO classification but works well for CON classification. However, adding both the attention gate and capsule module to the baseline model results in improvements of 1.43 and 1.5 for GGO and CON classifications, respectively, suggesting that the interaction between the two modules complements each other’s weaknesses, leading to better multi-class feature recognition. When compared with the complete architecture, disabling the capsule module leads to F1 score reductions of 3.2 and 3.09 for GGO and CON classifications, respectively, emphasizing the capsule’s role in capturing hierarchical relationships between categories. The effects of the attention gate and dual decoders on multi-class segmentation are also evaluated. Disabling the attention gate results in F1 score reductions of 2.62 and 3.24 for GGO and CON classifications, while disabling the dual decoders leads to F1 score reductions of 0.87 and 3.72 for the same classifications. These findings indicate that both the attention gate and dual decoders play significant roles in enhancing the performance of the model in multi-class segmentation tasks.
@@
In summary, the ablation experiments demonstrate that the capsule module, dual decoders, and attention gate are indispensable components of CAD-Unet, collectively contributing to the model’s outstanding performance in COVID-19 lung infections segmentation tasks. These results validate the effectiveness of our proposed architecture and provide valuable insights for future research in deep learning for medical image segmentation.
@@ 5 Discussion
5.1 Qualitative analysis
To visually demonstrate the efficacy of our proposed model in addressing both binary and multi-class segmentation tasks for COVID-19 lung CT lesion identification, we conduct a qualitative examination of the segmentation outcomes for both scenarios. Fig.5 provides a visual representation of partially predicted masks alongside the ground truth (GT) for the binary segmentation task, derived from three distinct datasets. The proposed model is compared with Unet++, nCoVSeg, SCOAT-Net, and PDAtt-Unet models. The results clearly demonstrate the superior performance of the predicted masks by our model, surpassing the other four models in terms of accurately encompassing the entire lesion region and exhibiting more precise and well-defined segmentation boundaries.
This demonstrates the capability of the model in effectively addressing binary segmentation challenges associated with COVID-19 lung lesions. The predicted masks produce by our model consistently demonstrate a propensity to encapsulate intricate lesion shapes while displaying sharper and more precise boundaries compared to the alternative models.
Fig.6 illustrates the visualization of partially predicted masks and GT for the multi-class segmentation task from two datasets, using our proposed model along with Unet++, SCOAT-Net, Anamnet, and D-TrAttUnet models. Comparing the predicted masks and GT of other models, the first row of images shows that the proposed model captures more complete CON class masks. The second row demonstrates the superior segmentation performance of our model on small target regions. The third row indicates that our model obtains more accurate target shape information, while the fourth row shows that we achieve more complete target segmentation masks in the GGO class.
@@ 5.1 Qualitative analysis
Model Computational cost Number of parameters Inference time
Unet 474.48 GFLOPs 7.96 M 2.7 ms
UNet++ 1698.59 GFLOPs 36.62 M 3.3 ms
Att-Unet 211.46 GFLOPs 10.29 M 4.3 ms
InfNet 90.69 GFLOPs 31.07 M 16.5 ms
AnamNet 311.57 GFLOPs 4.63 M 3.9 ms
SCOATNet 481.11 GFLOPs 10.21 M 7.7 ms
nCoVSegNet 636.72 GFLOPs 168.17 M 13.3 ms
PDEAtt-Unet 331.26 GFLOPs 14.44 M 8.9 ms
D-TrAttUnet 661.54 GFLOPs 104.16 M 13.0 ms
CAD-Unet 362.91 GFLOPs 15.04 M 23.4 ms
Table 12: Comparison of computational cost, model parameters, and inference time (Batch Size = 16)
5.2 Computational efficiency analysis
Efficient model design is critical for real-world clinical applications of medical image segmentation. Table 12 presents the computational cost, parameter count, and inference time for our proposed model and several baseline methods, where computations are performed with a batch size of 16 images.
Compared to the Transformer-based D-TrAttUnet (661.54 GFLOPs, 104.16M parameters), our Capsule Network-based model achieves competitive segmentation performance with significantly lower computational cost (362.91 GFLOPs) and model size (15.04M parameters). This demonstrates that capsule networks can achieve efficient feature representation with relatively low computational cost and parameter count, making them a promising alternative to Transformer-based architectures for medical image segmentation.
In terms of inference speed, our model achieves a batch inference time of 23.4 ms for 16 images, corresponding to an average of 1.4 ms per image, which is suitable for real-time medical diagnosis. However, compared to other models such as UNet (2.698 ms), UNet++ (3.300 ms), and Att-UNet (4.312 ms), our inference time is relatively longer. The increased latency may be attributed to the iterative routing mechanisms in capsule networks, which enhance feature representation but introduce additional computational overhead. Nevertheless, given its strong segmentation performance and efficient feature encoding, our model remains a viable choice for medical image analysis tasks that require both accuracy and robustness.
@@ 6 Conclusions
The outbreak of COVID-19 has presented significant challenges to global health. Lung CT imaging plays a crucial role in the early diagnosis and treatment of COVID-19 patients. This paper aims to develop an efficient and accurate lung CT lesion segmentation method to assist clinicians in quickly obtaining lesion information for more precise diagnosis and treatment. Our proposed CAD-Unet architecture combines the Unet encoder pathway and capsule encoder pathway in parallel to extract richer local, global, and long-range dependent features. Furthermore, to address the limitations of the BCE loss function in segmenting infection edges, we incorporate an edge loss component.
To assess the effectiveness of the proposed method, we conduct evaluations on two tasks: binary segmentation and multi-class segmentation. In addition, we compare our method with three baselines and five state-of-the-art architectures for COVID-19 segmentation. The experimental results demonstrate that our method outperforms all the mentioned architectures in common evaluation metrics for both binary and multi-class segmentation tasks. Furthermore, we perform ablation experiments to validate the effectiveness of the components in our proposed model. In future work, we will explore the potential of combining capsule networks with traditional segmentation models and applying them to a wider range of medical image segmentation tasks.
Acknowledgments
The authors would like to thank the Editor and the anonymous reviewers for their constructive comments to improve this manuscript. This work was supported by the National Natural Science Foundation of China (Grant Nos. 62463025, 62462051), by the Natural Science Foundation of Ningxia (Grant No. 2024AAC03038), and by the Youth Top Talent Training Program of Ningxia 2022 (Grant No. 2022012).
Appendix.
In the appendix, we attach the learning curves for binary and multi-class segmentation tasks in Fig.7 and Fig.8. The curves show the changes in loss, IoU, and F1 score during training, which help evaluate the model’s performance.
Fig. 7: Learning curves (1)
Fig. 8: Learning curves (2)